Jim saw a decrease in the quantity demanded for his firm's product from 8000 to 6000 units a week when he raised the price of the product from $200 to $250 Based on this information, the price elasticity of demand for Jim's product 

  • a. 1

  • b. less than 1

  • c. greater than 1

  • d. 0
Answer #1

ANSWER:

Price elasticity of demand = % change in quantity / % change in price

% change in quantity = (q2 - q1) / (q2 + q1) / 2 = (6000 - 8000) / (6000 + 8000) / 2 = -2000 / 7000 = -2 / 7

% change in price = (p2 - p1) / (p2 + p1) / 2 = (250 - 200) / (250 + 200) / 2 = 50 / 225 = 2 / 9

Price elasticity of demand = -2 / 7 / 2 / 9 = 9 / 7 = 1.28

so the correct answer is option c that is greater than 1.